OpenStudy (anonymous):

Use algebra to find all values of x between 0 and 2π that satisfy this equation. 4sin^2x-3=0

OpenStudy (solomonzelman):

4sin²x-3=0 4sin²x=3 sin²x=3/4 sinx=±√3/2

OpenStudy (solomonzelman):

So, 60º is good. (TO prove this, you can draw 30-60-90 and find the sine of 60. IF YOU WANT)

OpenStudy (anonymous):

Hm, so in radians the answer would be pi/3 and 5pi/6?

OpenStudy (solomonzelman):

:)

OpenStudy (anonymous):

Alright, thank you :D